A mixture of cement, sand and gravel is in the ratio 1-2-4. Determine amounts of each needed for 1 cu. yd. of concrete. (33 1/3% voids are allowed in the sand and 45% voids allowed in the gravel).
Let x = cu. yds. of cement.
Let y = cu. yds. of sand.
Let z = cu. yds. of gravel.

Equations provided with problem:
1) x + 2/3 y + .55z = 1 (cu. yd.)
2)y = 2x, z = 4x (from ratios of mixture).
Placing 2) in 1).
x + 4x/3 + 2.2x = 1
3x + 12x + 6.6 = 1
21.6x = 1
x = 0.046
Not the answer with problem. Where am I correct? Thanks.





Answer by stanbon(75887)

A mixture of cement, sand and gravel is in the ratio 1-2-4. Determine amounts of each needed for 1 cu. yd. of concrete. (33 1/3% voids are allowed in the sand and 45% voids allowed in the gravel).
-----
x + 2x + 4x = 1 cu yd
7x = 1
x = 1/7 = 0.1429
Let x = cu. yds. of cement = 0.1429 yd
Let y = cu. yds. of sand. = 2x = 0.2857 cu yd
Let z = cu. yds. of gravel = 4x = 0.5716 cu. yd
-----
Cheers,
Stan H.